OpenStudy (anonymous):

simplify the following expression (27/125)^2/3

OpenStudy (anonymous):

\[x^{\frac{ 2 }{ 3 }}=(\sqrt[3]{x})^2\]

OpenStudy (anonymous):

(27/125)^(2/3) =[(3/5)^3]^(2/3) =(3/5)^(2*3/3) =(3/5)^2 =9/25 =0.36

OpenStudy (anonymous):

so which ones right

OpenStudy (anonymous):

Azteck gives you the formula and I have given the solution totally. Now you decide.